# pasta: Pattern Analysis for Spatial Omics Data

### v2

@@ Abstract
Spatial omics assays allow for the molecular characterisation of cells in their spatial context. Notably, the two main technological streams, imaging-based and high-throughput sequencing-based, can give rise to very different data modalities. The characteristics of the two data types are well known in adjacent fields such as spatial statistics as point patterns and lattice data, and there is a wide range of tools available. This paper discusses the application of spatial statistics to spatially-resolved omics data and in particular, discusses various advantages, challenges, and nuances. This work is accompanied by a vignette, pasta (https://robinsonlabuzh.github.io/pasta), that showcases the usefulness of spatial statistics in biology using several R and Python packages.
@@ Introduction
Molecular profiling of cells in organs or tissues can be accomplished either in bulk or at the single-cell level. However, tissue must be dissociated, which may select against certain cell types and results in the loss of the spatial organization of the cells. This limitation is overcome with spatial profiling techniques, which can range from spatial proteomics (e.g., IMC, MIBI-TOF, 4i, CODEX, mIF [giesenHighlyMultiplexedImaging2014, kerenMIBITOFMultiplexedImaging2019, gutMultiplexedProteinMaps2018, blackCODEXMultiplexedTissue2021, tsujikawaQuantitativeMultiplexImmunohistochemistry2017]), to spatial transcriptomics assays based on fluorescence in situ hybridisation (e.g., FISH, MERFISH, CosMx, Xenium [langer-saferImmunologicalMethodMapping1982, chenSpatiallyResolvedHighly2015, janesickHighResolutionMapping2023]) and based on sequencing (e.g., 10x Visium and Visium HD, Slide-seq V1&V2 [stahlVisualizationAnalysisGene2016, oliveiraCharacterizationImmuneCell2024, rodriquesSlideseqScalableTechnology2019, stickelsHighlySensitiveSpatial2021]) to even spatial epigenomics [luSpatiallyResolvedEpigenomic2022]. Moreover, the combination of individual methods enabled spatial multiomic research [vandereykenMethodsApplicationsSinglecell2023]. The technologies and their application in biological research have been the topic of various detailed reviews [aspSpatiallyResolvedTranscriptomes2020, moffittEmergingLandscapeSpatial2022, raoExploringTissueArchitecture2021, mosesMuseumSpatialTranscriptomics2022, pallaSpatialComponentsMolecular2022, bressanDawnSpatialOmics2023].
@@
In this work, we focus on the concepts and application of exploratory statistical approaches for spatial data with a focus on downstream steps; our overview also centres around key differences between data modalities that result from current spatial profiling technologies. In this regard, we complement the ongoing discussion [zormpasMappingTranscriptomeRealizing2023, mosesVoyagerExploratorySinglecell2023] about the challenges in applying spatial statistics to spatial (transcript)omics data.
@@
Most spatial omics assays can be classified as either high-throughput sequencing (HTS) based or imaging-based. In HTS-based approaches, positional information is recorded using a predetermined array of spots or beads. Imaging-based approaches, however, either target the molecules of interest with fluorescent probes, ablate regions stained with a cocktail of antibodies via metal tag readouts, or amplify and sequence target sequences in situ. Several technologies are emerging, but the main trade-offs stem from the resolution, number of features, and sensitivity of the readout [raoExploringTissueArchitecture2021, parkSpatialOmicsTechnologies2022]. For example, HTS-based approaches capture the entire transcriptome (i.e., untargeted), but come at a resolution determined by the spot size. On the other hand, imaging-based approaches typically have lower depth of information (i.e., targeted) but show a higher resolution in comparison to HTS-based approaches. Note also that the landscape of spatially-resolved molecular technologies is rapidly changing, mostly in terms of higher resolution for HTS-based methods and comprehensive marker panels for imaging-based approaches (Figure 1, Technology) [moffittEmergingLandscapeSpatial2022, raoExploringTissueArchitecture2021, rademacherComparisonSpatialTranscriptomics2024].
@@
In terms of data analysis, the technological streams are quite distinct. HTS-based approaches collect data along regularly-spaced spots or beads resulting in a so-called regular lattice. In contrast, imaging-based approaches measure features at exact locations that can be assumed to have originated from a stochastic process known as a point process [raoExploringTissueArchitecture2021]. Thus, we will distinguish between lattice-based and point pattern-based spatial omic data. While the focus is often on the technology (c.f., imaging-based versus HTS-based), we argue that the distinction should lie at the data representation level [raoExploringTissueArchitecture2021]. This is important, since there are technologies that can be represented as both point patterns and lattice data, depending on the resolution and data processing. For example, imaging-based technologies that detect transcripts at subcellular resolution result in data where transcript locations can be represented as a point pattern, while for the same dataset, segmented cells can be represented either as an irregular lattice, or as a point pattern via their cell centroids. In general, if events of a point pattern are aggregated in specified regions (e.g., transcripts per cell from segmentation), we end up with lattice data (Figure 1, Data modality) [daleSpatialAnalysisGuide2014, pebesmaSpatialDataScience2023].
@@
Since we can represent some spatial omic datasets as either lattice or point pattern, it is important to further understand the assumptions and analysis strategies of such data modalities [cressieStatisticsSpatialData1993]. On the one hand, spatial point pattern analysis assumes that the point locations were generated by a stochastic process, a so-called event-based view of the data; the goal is to study the properties of this process [baddeleySpatialPointPatterns2015]. On the other hand, one could assume that the locations are fixed and known at the time of sampling and study the associated features at each location via an observation-based view of the data, while recognising that the observations in the lattice are not independent due to the spatial structure [anselinWhatSpecialSpatial1989].
Spatial data modalities have been present in other fields for decades, such as geography, and in particular, the field of spatial statistics offers a large analysis toolbox. For example, spatial omics data collected across a lattice can be analysed by exploratory tools, such as Local Moran’s and the bivariate Lee’s statistic. Imaging data generated by a stochastic process can be input for methods of point pattern analysis, including summaries such as Ripley’s or empty-space functions [zuurAnalysingEcologicalData2007, daleSpatialAnalysisGuide2014, baddeleySpatialPointPatterns2015, pebesmaSpatialDataScience2023] (Figure 1, Analysis).
There are already several tools for omics analysis that employ spatial statistics approaches. Amongst others, Voyager is a framework and a collection of use cases for lattice data, written in both R and Python [mosesVoyagerExploratorySinglecell2023]; it provides data structures and functionalities to compute spatial statistics in molecular biology. Voyager comes with comprehensive vignettes for several spatial profiling technologies. The methodological focus of their vignettes is on lattice data analysis and the spdep geospatial package. Since Voyager offers efficient implementations for lattice data analysis in R, we build on their framework in our vignettes for the lattice data component. Other methods written in R include: SPIAT, which contains various tools for spatial analysis, such as: neighbourhood analysis, local metrics and heterogeneity scores etc. [fengSpatialAnalysisSPIAT2023]; scFeatures, a toolbox comprising lattice data analysis functionality such as Moran’s and point pattern metrics as the -function [caoScFeaturesMultiviewRepresentations2022]; spicyR, which uses a compressed version of the global -function for cross-sample comparison [caneteSpicyRSpatialAnalysis2022]; lisaClust, a spatial domain detection method via LISA (local indicators of spatial association) curves [patrickSpatialAnalysisHighly2023]; mxfda, that uses point pattern summary functions for survival analysis [wrobelMxfdaComprehensiveToolkit2024]; spatialDM uses global and local bivariate Moran’s to score colocalisation of ligand-receptor pairs [liSpatialDMRapidIdentification2023]; and MERINGUE, which uses nearest-neighbour autocorrelation for spatially variable gene selection [millerCharacterizingSpatialGene2021]. Methods in Python include squidpy, a package with various spatial statistics tools for both point pattern and lattice data [pallaSquidpyScalableFramework2022] and PySAL a general Python library for spatial analysis [reyPySALPythonLibrary2010]. For a detailed view on general methods development in the field of spatial transcriptomics, we refer readers to the online Supplementary table of the museum of spatial transcriptomics paper [mosesMuseumSpatialTranscriptomics2022].
@@
Here, we will explore the application of spatial statistics to spatially-resolved omics data guided by the two main streams of data, point pattern-based and lattice-based methods. We will compare the two streams and show their strengths and limitations. Finally, we give an outlook on the challenges and research gaps in the field of spatial data analysis for omics data. Furthermore, we provide a vignette (https://robinsonlabuzh.github.io/pasta) that showcases analysis of data from multiple technologies, where concepts and assumptions are discussed in detail, with inline R and Python code. Overall, our resource Pattern Analysis for SpaTial omics datA (pasta) will highlight the usefulness and transferability of existing exploratory spatial statistics approaches in the context of spatial tissue profiling.
Results
Point Pattern Analysis
Definitions
Most imaging-based technologies give a high-resolution readout of either subcellular compartments (e.g., laser ablation pixels in IMC [giesenHighlyMultiplexedImaging2014]) or even individual transcripts (e.g., FISH [feminoVisualizationSingleRNA1998]). In molecule-level technologies, the expression of features (transcript quantifications, ion counts, fluorescent intensities) is recorded where they occur [aspSpatiallyResolvedTranscriptomes2020]. Often, we are interested in the distribution and composition of cell types. In order to annotate cells, we require (good) cell segmentation boundaries (an active area of research not discussed here; see [wangCellSegmentationImage2019, stringerCellposeGeneralistAlgorithm2021, greenwaldWholecellSegmentationTissue2022, petukhovCellSegmentationImagingbased2022, architSegmentAnythingMicroscopy2023]), allowing transcripts to be allocated to a given cell [fuBIDCellBiologicallyinformedSelfsupervised2024], and then cell type inference. Since the locations of the annotated cells are generated by a stochastic biological process, we can approximate e.g., the cell centroids as points and analyse their location as point patterns.
In the analysis of these point patterns, the goal is to make inferences on the point process that generated the data, not on the patterns themselves [baddeleySpatialPointPatterns2015, p. 127]. Points in a point pattern can also carry a mark, which is a (univariate or multivariate) variable associated with a point [baddeleySpatialPointPatterns2015, pp. 147, 563]. In cell biology, we could imagine both categorical (e.g., cell type or spatial domain labels) or continuous marks (e.g., gene expression).
@@ Results
When a point pattern with many cell types is considered, there are two approaches to formulate the setup. In the first variant, we consider cells in a tissue to depend on each other so their distribution is due to one overarching biological process; this is referred to as a multitype view. In contrast, when we consider the patterns to be created by point processes, we assume that these processes can be individually analysed; this represents the multivariable viewpoint (change in terminology in comparison to [baddeleySpatialPointPatterns2015]). The processes do not need to be independent, but the independence of components (and functions) is often used as a null model for multivariable analysis (see box Assumptions) [baddeleySpatialPointPatterns2015, pp. 565ff.].
Assumptions

### v3

# Harnessing the Potential of Spatial Statistics for Spatial Omics Data with pasta

@@ Abstract
Spatial omics allow for the molecular characterization of cells in their spatial context. Notably, the two main technological streams, imaging-based and high-throughput sequencing-based, give rise to very different data modalities. The characteristics of the two data types are well known in spatial statistics as point patterns and lattice data. In this perspective, we show the versatility of spatial statistics to quantify biological phenomena from local gene expression to tissue organization. As an example, we describe how to use exploratory metrics to address scientific questions in breast cancer, including cellular co-localization and gene co-expression analysis. We discuss technical concepts like window sampling, homogeneity and weight matrix construction and show their importance. We also provide pasta (https://robinsonlabuzh.github.io/pasta), an extensive analysis vignette for spatial statistics both using R and Python packages with further biology-driven applications.
@@ Introduction
Molecular profiling of cells in bulk or at the single-cell level can be accomplished by dissociating organs or tissues. Dissociation may select against certain cell types [denisenkoSystematicAssessmentTissue2020, dingSystematicComparisonSinglecell2020] and results in the loss of the spatial organization. This limitation is attenuated with spatial profiling techniques, which preserve the cellular spatial organization and range from spatial proteomics [giesenHighlyMultiplexedImaging2014, kerenMIBITOFMultiplexedImaging2019, gutMultiplexedProteinMaps2018, blackCODEXMultiplexedTissue2021, tsujikawaQuantitativeMultiplexImmunohistochemistry2017], to spatial transcriptomics based on fluorescent in situ hybridisation [langer-saferImmunologicalMethodMapping1982, chenSpatiallyResolvedHighly2015, janesickHighResolutionMapping2023] or on sequencing [stahlVisualizationAnalysisGene2016, oliveiraCharacterizationImmuneCell2024, rodriquesSlideseqScalableTechnology2019, stickelsHighlySensitiveSpatial2021, chenSpatiotemporalTranscriptomicAtlas2022], to spatial epigenomics [luSpatiallyResolvedEpigenomic2022] and spatial metabolomics [angelMatrixAssistedLaserDesorption2013, zhangMassSpectrometryImaging2023], or multi-omics [vandereykenMethodsApplicationsSinglecell2023]. The technologies and their application in biological research have been the topic of various detailed reviews [aspSpatiallyResolvedTranscriptomes2020, moffittEmergingLandscapeSpatial2022, raoExploringTissueArchitecture2021, mosesMuseumSpatialTranscriptomics2022, pallaSpatialComponentsMolecular2022, bressanDawnSpatialOmics2023].
@@
Spatial data modalities have been present in other fields for decades. They originated in mining and forestry and have expanded to other fields such as geography or ecology, and have been approached by spatial statistics methods [zuurAnalysingEcologicalData2007, daleSpatialAnalysisGuide2014, baddeleySpatialPointPatterns2015, pebesmaSpatialDataScience2023]. In this work, we focus on the concepts and application of exploratory statistical approaches for molecular spatial data; our overview also centres on key differences between data modalities that result from current spatial profiling technologies and the different types of questions that can be answered with spatial statistics. In this regard, we complement the ongoing discussion on the challenges in applying spatial statistics to spatial (transcript)omics data [zormpasMappingTranscriptomeRealizing2023, mosesVoyagerExploratorySinglecell2023, gonzalez2025analyzing].
Data modality is relevant to spatial data analysis
@@
Most spatial omics assays can be classified as either high-throughput sequencing (HTS)-based or imaging-based. In HTS-based approaches, positional information is recorded by barcoding a predetermined array of spots or beads. Imaging-based approaches, however, either target the molecules of interest with fluorescent probes, ablate regions stained with a cocktail of antibodies via metal tag readouts, or amplify and sequence target sequences in situ. Several technologies are emerging, but the main trade-offs stem from the resolution, number of features, and sensitivity of the readout [raoExploringTissueArchitecture2021, parkSpatialOmicsTechnologies2022]. For example, HTS-based approaches aim to profile the entire transcriptome (i.e., untargeted), but come at a resolution determined by the spot size. On the other hand, imaging-based approaches typically describe a lower number of features (i.e., targeted) but show a higher resolution. Note also that the landscape of spatially-resolved molecular technologies is rapidly changing, mostly in the direction of higher resolution for HTS-based methods and higher-plex marker panels for imaging-based approaches (Figure 1A; Supplementary Table S1) [moffittEmergingLandscapeSpatial2022, raoExploringTissueArchitecture2021, rademacherComparisonSpatialTranscriptomics2024].
@@
In terms of data analysis, HTS- and imaging-based outputs are quite distinct. HTS-based approaches collect data along regularly spaced spots or beads resulting in a so-called regular lattice. In contrast, imaging-based approaches measure features at exact locations that can be assumed to have originated from a random (stochastic)process known as a point process [raoExploringTissueArchitecture2021], leading to so-called point patterns. Although a random process might not be the correct assumption in biological development, the representation of point patterns can still be useful when correctly interpreted as we will show later. Thus, we will distinguish between lattice-based and point pattern-based spatial omics data. While the classification is often on the technology (c.f., imaging-based versus HTS-based), we argue that the distinction should lie at the data representation level [raoExploringTissueArchitecture2021]. This is important, since there are technologies that can be represented as both point patterns and lattice data depending on the resolution and data processing. For example, imaging-based data can be processed to describe the location of single transcripts at subcellular resolution (point pattern) while also identifying individual cells as either an irregular lattice when using the cell segmentation as the lattice structure or as a point pattern when considering cell centroids. In general, if events of a point pattern are aggregated into specified regions (e.g., transcripts per cell from segmentation), we end up with lattice data (Figure 1B) [daleSpatialAnalysisGuide2014, pebesmaSpatialDataScience2023, cressieStatisticsSpatialData1993].
Lattice and point pattern analysis represent two overlapping streams of spatial statistics
@@
Since we can represent some spatial omics datasets as either lattice or point pattern, it is important to further understand the assumptions underlying common data analysis strategies. On the one hand, point pattern analysis assumes that the point locations were generated by a random process, where each point is an event itself, a so-called event-based view of the data; its goal is to study the data generating process [baddeleySpatialPointPatterns2015].For example, if we are interested in the distribution of one cell type in an image, we would use point pattern methods to quantify this arrangement. On the other hand, one could assume that the locations are fixed and known at the time of sampling and study the associated features at each location via an observation-based view of the data, while recognizing that the observations in the lattice are not independent measurements due to their spatial dependence [anselinWhatSpecialSpatial1989]. For example, in spot-based technologies, locations are predefined. If we want to study a gene expression pattern in such an assay, we should account for the spatial dependence between spots (Figure 1C).
In addition, characterizing the marks (variable) of interest is also important. We consider two broad types of marks associated with spatial locations: categorical (qualitative, e.g. cell types or spatial domains) and numerical (quantitative, e.g. gene expression). In terms of dimensionality, marks can be analyzed one- (univariate), two- (bivariate) or three or more at a time (multivariate analysis). Figure 1C summarizes options to analyze each combination of the type of mark and its dimensionality. Defining characteristics of marks is relevant to choose appropriate metrics for the biological question in mind; e.g., infiltration of cell type A into a region dominated by cell type B is a question involving a categorical bivariate measure (two cell types involved) and can be analyzed as a type of co-localization analysis between type A and B cells.
Notably, there are more methods for the analysis of categorical marks with point patterns. For numerical marks, more options exist in lattice data analysis (Figure 1). This should serve as an indication of which stream can be useful, if the data modality from the experimental assay allows for both point pattern and lattice data analysis.
![[Uncaptioned image]](x1.png)
(Figure legend on next page.)
@@
Here, we will explore the application of spatial statistics to spatially-resolved omics data guided by concrete biological questions (Figure 1). We will further focus on the two main streams of data analysis, point pattern-based and lattice-based methods, describing their strengths and limitations. Finally, we will give an outlook on the challenges and research gaps in the field of spatial data analysis for omics data.
Biological applications of spatial statistics for spatial omics
Spatial autocorrelation reveals triple positive regions in breast tumours
We re-analyzed an imaging-based human breast cancer dataset [janesickHighResolutionMapping2023], where the cell-level data can be considered as both a point pattern or lattice data (Figure 1). Figure 2A shows the H&E image of the breast cancer section.
As the first feature of interest, we consider the measured gene expression, which is a numerical, continuous (cell-level) “mark”. As we are interested in relationships within the local neighbourhood of the cells rather than effects at specific locations or distances, lattice data is the appropriate analysis method (Figure 1).
In their publication, Janesick et al. [janesickHighResolutionMapping2023] investigate the expression of three clinically relevant hormone receptor genes: Estrogen receptor (ESR1), progesterone receptor (PGR) and human epidermal growth factor receptor 2 (ERBB2). The presence or absence of these receptors (indicated with +/-) can inform treatment strategies [janesickHighResolutionMapping2023]. The expression pattern of each receptor is quite distinct, with ERBB2 being more highly expressed compared to both ESR1 and PGR (Figure S1A-C). Lattice data analysis enables us to use a locally-scaled measure of expression to distinguish high and low expression regions. Since we are interested in localized hormone receptor-positive regions, we define the neighbourhood as the six nearest neighbours around each cell.
Univariate spatial autocorrelation allows us to identify regions where the expression of the hormone receptor genes is correlated among cells in their neighbourhood. Since autocorrelation measures such as local Moran’s [moranNotesContinuousStochastic1950, anselinLocalIndicatorsSpatial1995] are non-directional (e.g., local autocorrelation can be both high in regions where expression among neighbours is high or low as the focus in on similarity; for more details, see section “Spatial autocorrelation is a measure of similarity”), we rely on Moran’s scatter plot to identify the type of interaction [anselin2019moran]. In regions with “high-high” interaction, the expression among neighbours is both similar and high, which can be interpreted as a region positive for one of the receptors (Supplementary Figure S1D-F). Figure 2B shows the union of the individual positive univariate classifications of the three receptor genes and indicates single, double and triple positive receptor regions. One DCIS region (indicated by arrow) is lined with predominantly triple positive cells, but some local heterogeneity with double and single positive cells is apparent.
We also applied bi- or multivariate spatial correlation measures to classify the different regions. Similar to spatial autocorrelation, bivariate autocorrelation measures the similarity of expression of two genes in a local neighbourhood. Supplementary Figure S2A-C show bivariate local Lee’s [leeDevelopingBivariateSpatial2001] correlation measures for the pairs of receptors ERBB2-ESR1, ERBB2-PGR and ESR1-PGR, respectively; regions with high values indicate regions where expression of both receptor genes is similar (“high-high” or “low-low”). The different scales of the local Lee’s L values are the result of the different expression scales (c.f., Supplementary Figure S1A-C). A comparison of gene expression values and the local Lee’s value allows us to identify double positive regions.
Multivariate spatial correlation metrics extend the concept of spatial correlation to more than two genes. The main difficulty is finding an appropriate metric to measure similarity of more than two marks. In multivariate Geary’s , the respective metric is the weighted average of the squared differences of the expression vectors of the three genes. Supplementary Figure S2D-F shows an analysis of multivariate local Geary’s [anselinLocalIndicatorMultivariate2019] for the three hormone receptors. Cells with a high local Geary’s value exhibit a gene expression profile that is highly dissimilar (squared difference is high) to their local neighbours and indicates local heterogeneity. Interestingly, the triple positive region identified using Moran’s scatter plot shows high multivariate Geary’s values. At first, this seems counterfactual to what we have identified using the Moran’s scatter plot. A closer look reveals that the region is very heterogeneous with a lot of single and double-positive regions with varying scales of expression of the individual genes. Generally, there is a difficulty in interpreting multivariate autocorrelation measures as they are not simply the sum of the respective univariate counterparts and should therefore be mainly used to identify interesting regions (c.f., Anselin et al. [anselinLocalIndicatorMultivariate2019]).
In summary, we have investigated the expression patterns of the three clinically-relevant receptor genes, quantifying single-, double- and triple-positive receptor regions in the tumour using lattice data analysis. Spatial statistics was able to recapitulate the original findings in Janesick et al. [janesickHighResolutionMapping2023] (Figure 5), and in addition, spatial exploratory statistics provide a quantification of the strength of the various spatial correlations as well as an indication of the significance of the effects.
Point pattern analysis quantifies different degrees of tumour invasion into ductal carcinomas
Another analysis to consider is the exploration of the spatial distribution of cell types in the Janesick et al. [janesickHighResolutionMapping2023] dataset. Since cell types are categorical features, point pattern analysis is a powerful choice [baddeleySpatialPointPatterns2015] (Figure 1). Here we consider the DCIS 1, DCIS 2 and invasive tumour cell types and aggregate all other cell types into the “Other” category (Figure 3A). In this example, we calculate Besag’s cross function to quantify the co-localization of the three cell types in space [besag1977contribution, ripleySecondOrderAnalysisStationary1976]. Note that the observation window was restricted to the region where most of the cells were found to account for the underlying inhomogeneous distribution of cells (c.f., Figure 3). There are other options to account for this inhomogeneity, which are compared in Figure S3. In general, Besag’s calculates the average number of neighbours of a certain cell type as a function of the radius. This can used to estimate the magnitude of co-occurrence relative to a scenario where all cells are randomly distributed in the observation window. In Besag’s , this corresponds to a value of 0 across all radii. Positive deviation indicates more cells are observed than expected by chance, implying clustering; negative deviation indicates less cells observed than expected by chance, implying spacing of the cells.
First, we consider univariate co-localization, i.e., whether cells are found more often than expected in the vicinity of other cells from the same type. Figure 3B shows Besag’s curve plots for each pair of cell types. We note that all cell types exhibit curves above the (dashed black) reference line, which means that we observe, on average, more cells of the same type within a certain radius than expected by chance. This represents a higher degree of clustering than a fully random, or completely spatial random (CSR), point pattern (see Technical description below; c.f., box “Importance of scale and uniformity in point pattern analysis”). Thus, all three cell types have a propensity higher than expected by chance to occur in the neighbourhood of their own type (Figure 3A) and this clustering is strong across all scales (or radii, ).
Besag’s can quantify bivariate interactions, i.e., the co-localization of a pair of cell types, as shown in Figure 3B. We note that the combination of DCIS 1 and DCIS 2 is below the dashed CSR line. This means that we observe less cells of the other type within a specific radius than expected by chance. This can be interpreted such that there is, on average, spacing between the respective cell types. Of interest is the co-localization of the two DCIS subtypes and invasive tumour cells: DCIS 1 cells and invasive tumour cells are strongly spaced, thus clearly co-occur less frequently than expected at random. On the other hand, DCIS 2 cells and invasive tumour cells are close to the CRS line which means they occur as often as expected by chance, an indication of mixing of the two cells types in space.
In summary, we have recapitulated the main findings on DCIS cells presented in Janesick et al., Figure 4, using point pattern analysis [janesickHighResolutionMapping2023]. DCIS subtypes show a different degree of co-localization with invasive tumour cells, with DCIS 1 regions showing no co-localization with invasive tumour while DCIS 2 exhibits a degree of co-localization with invasive tumour cells that can be expected by chance. Using spatial point pattern analysis, we have systematically quantified the co-localization of three cell types in breast cancer.
Further analyses using spatial statistics
In addition to the analyses described above, spatial statistics could be used for various other analyses.
Univariate global spatial autocorrelation methods such as Moran’s can be used to identify the strength of spatial correlation across all genes in the panel. This information can be useful to compare spatial variability of different genes within one sample and the specified neighbourhood. In addition, by constructing a spatial weight matrix based on contiguous neighbours (i.e., cells in physical contact with each other), we can identify pairs of ligand-receptor genes via quantification of their bivariate spatial correlation [liSpatialDMRapidIdentification2023]. Alternatively, adjusting the weight matrix to neighbours based on a certain distance, we can investigate paracrine signalling up to the distance specified.
Investigating the spatial arrangement of cells with point pattern analysis, we could further use mark correlation functions to quantify the scale of spatial correlation of the expression of a gene of interest. In addition, we could profile the co-localization of cell types using nearest-neighbour approaches such as the nearest-neighbour distance and orientation. These metrics would provide us with a more quantitative readout of the spatial organization than we can identify by eye. “Dot” (or “-to-any”) functions would allow us to make statements about how one cell type is localized with respect to a set of other cell types.
Moreover, we could also focus our analyses to regions of interest such as the single-, double- and triple-positive tumour regions to make statements about localized gene expression and the arrangement of cell types within these regions. The observation scale is very important in spatial analysis as it will impact the interpretation of spatial relationship, as we will further explain below.
![[Uncaptioned image]](x3.png)
(Figure legend on next page.)
Technical considerations
Inference about the point process
@@
In the analysis of point patterns, the goal is to make inferences on the point process that generated the data, not on the patterns themselves [baddeleySpatialPointPatterns2015, p. 127]. When considering a point pattern with many cell types, there are two possible formulations. In the first variant, we assume that cells in a section depend on each other so their distribution is due to one overarching biological process; this is referred to as a multitype view. In contrast, when we consider the patterns are created by point processes, we assume that these processes can be individually analyzed; this represents the multivariable viewpoint (change in terminology in comparison to [baddeleySpatialPointPatterns2015]). Generally, it is not easy to define whether the entire pattern of several cell types should be represented as a single point process (multitype analysis). Notably, analyzing cell types individually and considering them as separate point processes might underestimate dependencies among the cells (multivariable analysis) [baddeleySpatialPointPatterns2015, pp. 565 ff.].
Point pattern analysis quantifies spatial interactions across scales
A central idea in point pattern analysis is that of an -neighbourhood. The -neighbourhood describes the spatial dependence in a disk with radius . In such a neighbourhood, different aspects can be quantified, leading to different dependent variables per function. The expected number without any spatial interaction (CSR or Poisson process) can be calculated in the -neighbourhood. Positive deviations from the CSR value indicate clustering whereas negative deviations indicate spacing [baddeleySpatialPointPatterns2015, p. 204]. Therefore, point pattern metrics are functions that quantify spatial dependence across scales.
The two most well-known classes in point pattern analysis are correlation and spacing functions. The two approaches are highly complementary and there is a duality between them [baddeleySpatialPointPatterns2015, p. 255]. Correlation is usually assessed with Ripley’s function or its variance-stabilized version, Besag’s . Here, the number of points falling in an -neighbourhood is quantified. These raw counts are corrected for varying intensities as well as edge effects (see Box Importance of scale and uniformity in point pattern analysis) [ripleyFindingEdgePoisson1977, besag1977contribution][baddeleySpatialPointPatterns2015, p. 203 ff.]. Spacing can be assessed with the nearest-neighbour function . Here, the probability of finding the nearest-neighbour of a point in an -neighbourhood is calculated [baddeleySpatialPointPatterns2015, p. 261 ff.]. Both spacing and correlation quantify different aspects of a point pattern and a spatial analysis should consider both complementary options [baddeleySpatialPointPatterns2015, p. 295 ].
Window sampling and small world impact representative sample collection